OpenStudy (anonymous):

Mr.Volley's car lot sells an average of 15 cars and 6 trucks per week how many cehiles would this be for an 8-week period please explain

OpenStudy (anonymous):

15+6=21 21*8weeks = 168 vehicles

OpenStudy (anonymous):

to find the total number of vehicles you must add the number of cars and the number of trucks together. This is the amount of vehicles per week, then you just multiply the number of vehicles by the number of weeks which is 21*8=168 vehicles.